# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 6.1. Lifting operators and added mass
$C_{2}Q$ $=\frac{2\xi}{3}\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}(\Delta Q-Q)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times(\Delta Q-Q)n\,\mathrm{d}\Gamma\end{pmatrix},\enspace C_{3}\binom{\ell}{\omega}=-2\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}\mathbb{D}((\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega))n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times\mathbb{D}((\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega))n\,\mathrm{d}\Gamma\end{pmatrix}.$
Let us note that $M$ above is generally referred to as the added mass operator. From, [16, Lemma 4.6], see also [18, Lemma 4.3], we recall that the matrix $K$ from (6.16) is invertible.
@@ 6.2. The fluid-structure operator
(6.17) $A_{\mathrm{fs}}(\widehat{Q})$ $\coloneqq\begin{pmatrix}A_{2}&\mathbb{P}\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})&-A_{2}\mathbb{P}D_{\mathrm{fl}}\\ -\tilde{S}_{\xi}(\widehat{Q})\nabla&\Delta_{\mathrm{N}}^{1}&0\\ K^{-1}C_{1}&K^{-1}C_{2}&K^{-1}C_{3}\end{pmatrix}.$
@@
Let $(f_{1},f_{2},f_{3},f_{4})\in\mathrm{X}$ and $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$. Then $(v,Q,\ell,\omega,p)$ with $v\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}$, $Q\in\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})$, $(\ell,\omega)\in\mathbb{C}^{3}\times\mathbb{C}^{3}$ and $p\in\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ is a solution to (6.15) if and only if
$(\lambda\operatorname{Id}-A_{\mathrm{fs}}(\widehat{Q}))\begin{pmatrix}\mathbb{P}v\\ Q\\ \ell\\ \omega\end{pmatrix}&=\begin{pmatrix}\mathbb{P}f_{1}\\ f_{2}\\ K^{-1}f_{3}\\ K^{-1}f_{4}\end{pmatrix},\enspace(\operatorname{Id}_{3}-\mathbb{P})v=(\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega)\enspace\text{and}\enspace\\ p&=N(\Delta\mathbb{P}v\cdot n)-\lambda N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n).\end{aligned}\right.$
@@
Let $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, and recall the fluid-structure operator $A_{\mathrm{fs}}(\widehat{Q})$ from (6.17). Then there exists $\mu\geq 0$ such that $-A_{\mathrm{fs}}(\widehat{Q})+\mu$ is sectorial on $\mathrm{X}$ with spectral angle $\phi_{-A_{\mathrm{fs}}(\widehat{Q})+\mu}<\nicefrac{{\pi}}{{2}}$. In particular, $-A_{\mathrm{fs}}(\widehat{Q})+\mu$ has maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}$.
Proof.
First, we decompose the fluid-structure operator $A_{\mathrm{fs}}(\widehat{Q})$ into $A_{\mathrm{fs}}(\widehat{Q})\coloneqq A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})+B_{\mathrm{fs}}$, where
$A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})\coloneqq\begin{pmatrix}-A_{\xi}(\widehat{Q})&0\\ 0&0\end{pmatrix}\enspace\text{and}\enspace B_{\mathrm{fs}}\coloneqq\begin{pmatrix}0&B_{1}\\ B_{2}&B_{3}\end{pmatrix}\coloneqq\begin{pmatrix}0&0&-A_{2}\mathbb{P}D_{\mathrm{fl}}\\ 0&0&0\\ K^{-1}C_{1}&K^{-1}C_{2}&K^{-1}C_{3}\end{pmatrix}$
for $A_{\xi}(\widehat{Q})$ from (6.8) and the Stokes operator $A_{2}$ on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$. Besides, the Dirichlet lifting operator $D_{\mathrm{fl}}$ was made precise in (6.11), and the operators $C_{1}$, $C_{2}$ and $C_{3}$ and the matrix $K$ are defined in (6.16).
The sectoriality of $-A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})$ on $\mathrm{X}$ with spectral angle less than $\nicefrac{{\pi}}{{2}}$ carries over from Section 6(a). For the sectoriality of $-A_{\mathrm{fs}}(\widehat{Q})$ up to a shift, we thus verify the relative $-A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})$-boundedness of $B_{\mathrm{fs}}$ with arbitrarily small $-A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})$-bound. To this end, consider $(v,Q,\ell,\omega)\in\mathrm{D}(A_{\mathrm{fs}})=\mathrm{D}(A_{\mathrm{fs},\mathrm{I}})$. Thanks to mapping properties of the Stokes operator as well as $D_{\mathrm{fl}}\in\mathcal{L}(\mathbb{C}^{6},\mathrm{H}^{2}(\mathcal{F}_{0}))$, we deduce that
@@
This shows that $B_{1}\in\mathcal{L}(\mathbb{C}^{3}\times\mathbb{C}^{3},\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3}))$, so $B_{1}$ is a bounded perturbation. Moreover, we recall from the proof of [26, Theorem 3.11] that the operators $C_{1}$ and $C_{3}$ are relatively bounded with relative bound zero. For $C_{2}$, we invoke classical properties of the trace, interpolation and Young’s inequality to deduce that for any $\delta>0$, and with $\varepsilon>0$ small and $\theta\in(0,1)$, it holds that
@@
Thus, $C_{2}$ is a relatively bounded perturbation with relative bound zero. In total, $B_{\mathrm{fs}}$ is a relatively bounded perturbation of $A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})$ with $A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})$-bound zero. Perturbation theory for sectorial operators, see e. g. [33, Cor. 3.1.6], then yields the existence of $\mu\geq 0$ such that $-A_{\mathrm{fs}}(\widehat{Q})+\mu=-A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})-B_{\mathrm{fs}}+\mu$ is sectorial on $\mathrm{X}$ with spectral angle $\phi_{-A_{\mathrm{fs}}(\widehat{Q})+\mu}<\nicefrac{{\pi}}{{2}}$. The second part of the assertion follows by the known equivalence of sectoriality with angle $<\nicefrac{{\pi}}{{2}}$ and maximal $\mathrm{L}^{p}$-regularity on Hilbert spaces. ∎
@@
For simplicity, we write $A_{\mathrm{fs}}\coloneqq A_{\mathrm{fs}}(0)$ in the sequel. It can be shown that the shift $\mu\geq 0$ in Subsection 6.2 can be be chosen equal to $s(A_{\mathrm{fs}})\coloneqq\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A_{\mathrm{fs}})\}$ of $A_{\mathrm{fs}}$, see for instance [33, Cor. 3.5.3]. Hence, to remove the shift in Subsection 6.2, and thus to obtain the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of $A_{\mathrm{fs}}$ on $\mathrm{X}$, we need to prove that the spectral bound satisfies $s(A_{\mathrm{fs}})<0$.
@@
Let now $\lambda\in\mathbb{C}$ and $(v,Q,\ell,\omega,p)\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{C}^{3}\times\mathbb{C}^{3}\times\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ satisfy (6.15) with $f_{i}=0$ for $i=1,\dots,4$. Next, we test the equation satisfied by $v$ in (6.15) by $\overline{v}$, the equation for $Q$ by $\Delta\overline{Q}$, the equation for $\ell$ by $\overline{\ell}$ and the equation for $\omega$ by $\overline{\omega}$. Integrating by parts, invoking Section 6(b) to handle the liquid crystals part, observing that $S_{\xi}(0)(-\Delta+\operatorname{Id}_{3})Q=\nicefrac{{2\xi}}{{3}}(\Delta Q-Q)$, making use of the interface condition from (2.8), and recalling $\mathrm{T}(v,p,Q)$ from (6.4), we derive that
00 0 $=\lambda\left(\|v\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\|\nabla Q\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+m_{\mathcal{S}}|\ell|^{2}+J_{0}\omega\cdot\overline{\omega}\right)$
@@
It follows that $\operatorname{Re}\lambda\leq 0$. Moreover, if $\operatorname{Re}\lambda=0$, then $\mathbb{D}(v)=0$ and $Q=0$ on $\mathcal{F}_{0}$, so $v$ is especially constant. Due to $v=0$ on $\partial\mathcal{O}$, we infer that $v=0$ on $\mathcal{F}_{0}$. The interface condition $v=\ell+\omega\times y$ on $\partial\mathcal{S}_{0}$ then yields that $\ell=\omega=0$. From (6.15), it follows that $\nabla p=0$, and $p\in\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ rules out the constants, so $p=0$ and thus $(v,Q,\ell,\omega,p)=(0,0,0,0,0)$. The above arguments imply that $s(A_{\mathrm{fs}})<0$, resulting in the sectoriality and then also the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of $-A_{\mathrm{fs}}$ on $\mathrm{X}$ even without shift.
@@
By the sectoriality of $-A_{\mathrm{fs}}$ with spectral angle $\phi_{-A_{\mathrm{fs}}}<\nicefrac{{\pi}}{{2}}$, the operator $A_{\mathrm{fs}}$ generates a bounded analytic semigroup $\left(\mathrm{e}^{tA_{\mathrm{fs}}}\right)_{t\geq 0}$, see e. g. [33, Thm. 3.3.2]. Thanks to $s(A_{\mathrm{fs}})<0$, this semigroup is even of negative exponential type. Thus, we may choose $\eta_{0}=-s(A_{\mathrm{fs}})>0$ and consider the operator $A_{\mathrm{fs}}+\eta$ for $\eta\in[0,\eta_{0})$, still generating a bounded analytic semigroup of negative exponential type thanks to the fact that $s(A_{\mathrm{fs}}+\eta)<0$, to deduce the estimate with exponential weights. ∎
@@ 7. Local strong well-posedness for large data
Let $p>4$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, where $\mathrm{X}_{\gamma}$ has been introduced in (2.10), and consider $(f_{1},f_{2},f_{3},f_{4})=(0,0,0,0)$. Then for $\mu>\mu_{0}$, where $\mu_{0}\geq 0$ is as in Theorem 6.1(a), there is a unique solution $(z^{*},p^{*})\in\tilde{\mathbb{E}}_{\infty}$ to (6.5). In particular, we have $\|z^{*}-z_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\to 0$ as $T\to 0$.
For the reference solution $(z^{*},p^{*})$ from Section 7 and a solution $(z,p)$ to (5.11)–(5.13), we define $\widehat{v}\coloneqq v-v^{*}$, $\widehat{Q}\coloneqq Q-Q^{*}$, $\widehat{\ell}\coloneqq\ell-\ell^{*}$, $\widehat{\omega}\coloneqq\omega-\omega^{*}$ and $\widehat{p}\coloneqq p-p^{*}$. For $\mu>\mu_{0}$, with $\mu_{0}\geq 0$ from Theorem 6.1(a), the emerging $(\widehat{z},\widehat{p})=(\widehat{v},\widehat{Q},\widehat{\ell},\widehat{\omega},\widehat{p})$ solves
(7.1) $\partial_{t}{\widehat{v}}+(\mu-\Delta){\widehat{v}}+\nabla{\widehat{p}}-\operatorname{div}S_{\xi}(Q_{0})(\Delta-\operatorname{Id}_{3})\widehat{Q}&=F_{1}(\widehat{z},\widehat{p}),\enspace\operatorname{div}{\widehat{v}}=0,&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{\widehat{Q}}}+\tilde{S}_{\xi}(Q_{0})\nabla\widehat{v}+(\mu-(\Delta-\operatorname{Id}_{3})){\widehat{Q}}&=F_{2}(\widehat{z}),&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\widehat{\ell}})^{\prime}+\mu\widehat{\ell}+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(\widehat{v},\widehat{p},\widehat{Q})n\,\mathrm{d}\Gamma&=F_{3}(\widehat{z},\widehat{p}),&&\enspace\text{in}\enspace(0,T),\\ J_{0}({\widehat{\omega}})^{\prime}+\mu\widehat{\omega}+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(\widehat{v},\widehat{p},\widehat{Q})n\,\mathrm{d}\Gamma&=F_{4}(\widehat{z},\widehat{p}),&&\enspace\text{in}\enspace(0,T),\\ {\widehat{v}}={\widehat{\ell}}+{\widehat{\omega}}\times y,\enspace\text{on}\enspace(0,T)\times\partial\mathcal{S}_{0},\enspace{\widehat{v}}=0,&\enspace\text{on}\enspace(0,T)\times\partial\mathcal{O},\enspace\partial_{\nu}\widehat{Q}=0,&&\enspace\text{on}\enspace(0,T)\times\partial\mathcal{F}_{0},\\ \widehat{v}(0)=0,\enspace\widehat{Q}(0)=0,\enspace\widehat{\ell}(0)&=0\enspace\text{and}\enspace\widehat{\omega}(0)=0.\end{aligned}\right.$
@@
$\mathcal{T}(v,p,Q)\coloneqq\mathbb{O}(t)^{\top}\mathrm{T}(\mathbb{O}(t)v(t,y),p(t,y),Q(t,y))\mathbb{O}(t)$
for the rotation matrix $\mathbb{O}(t)$ related to (5.1), $\mathrm{T}(v,p,Q)$ as in (6.4) and $(z,p)=(\widehat{z}+z^{*},\widehat{p}+p^{*})$, we have
(7.2) $F_{1}(\widehat{z},\widehat{p})$ $\coloneqq(\mathcal{L}_{1}-\Delta)v+\mu v-\mathcal{M}v-\mathcal{N}(v)-(\mathcal{G}-\nabla)p+\mathcal{B}_{\tau_{\mathrm{l}}}(Q)$
@@
$F_{2}(\widehat{z})$ $\coloneqq(\mathcal{L}_{2}-\Delta)Q+\mu Q-(\partial_{t}Y\cdot\nabla)Q-(v\cdot\nabla)Q+\bigl{(}\mathcal{S}(v,Q)+\tilde{S}_{\xi}(Q_{0})\nabla v\bigr{)}$

### v2

@@ 6.1. Lifting operators and added mass
$C_{2}Q$ $=\frac{2\xi}{3}\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}(\Delta Q-Q)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times(\Delta Q-Q)n\,\mathrm{d}\Gamma\end{pmatrix},\kern 5.0ptC_{3}\binom{\ell}{\omega}=-2\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}\mathbb{D}((\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega))n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times\mathbb{D}((\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega))n\,\mathrm{d}\Gamma\end{pmatrix}$.
Let us note that $M$ above is generally referred to as the added mass operator. From, [17, Lemma 4.6], see also [19, Lemma 4.3], we recall that the matrix $K$ from (6.16) is invertible. The purpose of the added mass operator is that allows for a reformulation of the surface integrals without the pressure. More precisely, thanks to the invertibility of $K$, it is possible to reformulate the equations for the rigid body velocities including the surface integrals as a problem of $\mathbb{P}v$, $Q$ and a lifted problem involving the rigid body velocities $\ell$ and $\omega$ themselves.
@@ 6.2. The fluid-structure operator
[…]
@@
Proposition 6.6.
Let $(f_{1},f_{2},f_{3},f_{4})\in\mathrm{X}$ and $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$. Then $(v,Q,\ell,\omega,\pi)$ with $v\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}$, $Q\in\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})$, $(\ell,\omega)\in\mathbb{C}^{3}\times\mathbb{C}^{3}$ and $\pi\in\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ is a solution to (6.15) if and only if
$\left\{\begin{aligned} (\lambda\operatorname{Id}-A_{\mathrm{fs}}(\widehat{Q}))\begin{pmatrix}\mathbb{P}v\\ Q\\ \ell\\ \omega\end{pmatrix}&=\begin{pmatrix}\mathbb{P}f_{1}\\ f_{2}\\ K^{-1}f_{3}\\ K^{-1}f_{4}\end{pmatrix},\kern 5.0pt(\operatorname{Id}_{3}-\mathbb{P})v=(\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega)\kern 5.0pt\text{and}\kern 5.0pt\\ \pi&=N(\Delta\mathbb{P}v\cdot n)-\lambda N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n).\end{aligned}\right$.
Proposition 6.7.
@@
Proof.
@@
$A_{\mathrm{fs},\mathrm{I}}(\widehat{Q})\coloneqq\begin{pmatrix}-A_{\xi}(\widehat{Q})&0\\ 0&0\end{pmatrix}\kern 5.0pt\text{and}\kern 5.0ptB_{\mathrm{fs}}\coloneqq\begin{pmatrix}0&B_{1}\\ B_{2}&B_{3}\end{pmatrix}\coloneqq\begin{pmatrix}0&0&-A_{2}\mathbb{P}D_{\mathrm{fl}}\\ 0&0&0\\ K^{-1}C_{1}&K^{-1}C_{2}&K^{-1}C_{3}\end{pmatrix}$
$A_{\xi}(\widehat{Q}) fro$m (6.8) and the Stokes operator $A_{2}$ on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$. Besides, the Dirichlet lifting operator $D_{\mathrm{fl}}$ was made precise in (6.11), and the operators $C_{1}$, $C_{2}$ and $C_{3}$ and the matrix $K$ are defined in (6.16).
[…]
@@
This shows that $B_{1}\in\mathcal{L}(\mathbb{C}^{3}\times\mathbb{C}^{3},\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3}))$, so $B_{1}$ is a bounded perturbation. Moreover, we recall from the proof of [27, Theorem 3.11] that the operators $C_{1}$ and $C_{3}$ are relatively bounded with relative bound zero. For $C_{2}$, we invoke classical properties of the trace, interpolation and Young’s inequality to deduce that for any $\delta>0$, and with $\varepsilon>0$ small and $\theta\in(0,1)$, it holds that
@@
[…]
@@
For simplicity, we write $A_{\mathrm{fs}}\coloneqq A_{\mathrm{fs}}(0)$ in the sequel. It can be shown that the shift $\mu\geq 0$ in Subsection 6.2 can be be chosen equal to $s(A_{\mathrm{fs}})\coloneqq\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A_{\mathrm{fs}})\}$ of $A_{\mathrm{fs}}$, see for instance [35, Cor. 3.5.3]. Hence, to remove the shift in Subsection 6.2, and thus to obtain the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of $A_{\mathrm{fs}}$ on $\mathrm{X}$, we need to prove that the spectral bound satisfies $s(A_{\mathrm{fs}})<0$.
@@
Let now $\lambda\in\mathbb{C}$ and $(v,Q,\ell,\omega,\pi)\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{C}^{3}\times\mathbb{C}^{3}\times\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ satisfy (6.15) with $f_{i}=0$ for $i=1,\dots,4$. Next, we test the equation satisfied by $v$ in (6.15) by $\overline{v}$, the equation for $Q$ by $\Delta\overline{Q}$, the equation for $\ell$ by $\overline{\ell}$ and the equation for $\omega$ by $\overline{\omega}$. Integrating by parts, invoking Section 6(b) to handle the liquid crystals part, observing that $S_{\xi}(0)(-\Delta+\operatorname{Id}_{3})Q=\nicefrac{{2\xi}}{{3}}(\Delta Q-Q)$, making use of the interface condition from (2.8), and recalling $\mathrm{T}(v,\pi,Q)$ from (6.4), we derive that
$0$ $=\lambda\left(\|v\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+\|\nabla Q\|_{\mathrm{L}^{2}(\mathcal{F}_{0})}^{2}+m_{\mathcal{S}}|\ell|^{2}+J_{0}\omega\cdot\overline{\omega}\right)$
@@
It follows that $\operatorname{Re}\lambda\leq 0$. Moreover, if $\operatorname{Re}\lambda=0$, then $\mathbb{D}(v)=0$ and $Q=0$ on $\mathcal{F}_{0}$, so $v$ is especially constant. Due to $v=0$ on $\partial\mathcal{O}$, we infer that $v=0$ on $\mathcal{F}_{0}$. The interface condition $v=\ell+\omega\times y$ on $\partial\mathcal{S}_{0}$ then yields that $\ell=\omega=0$. From (6.15), it follows that $\nabla\pi=0$, and $\pi\in\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ rules out the constants, so $\pi=0$ and thus $(v,Q,\ell,\omega,\pi)=(0,0,0,0,0)$. The above arguments imply that $s(A_{\mathrm{fs}})<0$, resulting in the sectoriality and then also the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of $A_{\mathrm{fs}}$ on $\mathrm{X}$ even without shift.
@@
By the sectoriality of $A_{\mathrm{fs}}$ with spectral angle $\phi_{-A_{\mathrm{fs}}}<\nicefrac{{\pi}}{{2}}$, the operator $A_{\mathrm{fs}}$ generates a bounded analytic semigroup $\left(\mathrm{e}^{tA_{\mathrm{fs}}}\right)_{t\geq 0}$, see e. g. [35, Thm. 3.3.2]. Thanks to $s(A_{\mathrm{fs}})<0$, this semigroup is even of negative exponential type. Thus, we may choose $\eta_{0}=-s(A_{\mathrm{fs}})>0$ and consider the operator $A_{\mathrm{fs}}+\eta$ for $\eta\in[0,\eta_{0})$, still generating a bounded analytic semigroup of negative exponential type thanks to the fact that $s(A_{\mathrm{fs}}+\eta)<0$, to deduce the estimate with exponential weights. ∎
@@ 7. Local strong well-posedness for large data
Proposition 7.1.
Let $p>4$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, where $\mathrm{X}_{\gamma}$ has been introduced in (2.10), and consider $(f_{1},f_{2},f_{3},f_{4})=(0,0,0,0)$. Then for $\mu>\mu_{0}$, where $\mu_{0}\geq 0$ is as in Theorem 6.1(a), there is a unique solution $(z^{*},\pi^{*})\in\tilde{\mathbb{E}}_{\infty}$ to (6.5). In particular, we have $\|z^{*}-z_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\to 0$ as $T\to 0$.
@@
(7.1) $\left\{\begin{aligned} \partial_{t}{\widehat{v}}+(\mu-\Delta){\widehat{v}}+\nabla{\widehat{\pi}}-\operatorname{div}S_{\xi}(Q_{0})(\Delta-\operatorname{Id}_{3})\widehat{Q}&=F_{1}(\widehat{z},\widehat{\pi}),\kern 5.0pt\operatorname{div}{\widehat{v}}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{\widehat{Q}}}+\tilde{S}_{\xi}(Q_{0})\nabla\widehat{v}+(\mu-(\Delta-\operatorname{Id}_{3})){\widehat{Q}}&=F_{2}(\widehat{z}),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\widehat{\ell}})^{\prime}+\mu\widehat{\ell}+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(\widehat{v},\widehat{\pi},\widehat{Q})n\,\mathrm{d}\Gamma&=F_{3}(\widehat{z},\widehat{\pi}),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}({\widehat{\omega}})^{\prime}+\mu\widehat{\omega}+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(\widehat{v},\widehat{\pi},\widehat{Q})n\,\mathrm{d}\Gamma&=F_{4}(\widehat{z},\widehat{\pi}),&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ {\widehat{v}}={\widehat{\ell}}+{\widehat{\omega}}\times y,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0},\kern 5.0pt{\widehat{v}}=0,&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\partial_{n}\widehat{Q}=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ \widehat{v}(0)=0,\kern 5.0pt\widehat{Q}(0)=0,\kern 5.0pt\widehat{\ell}(0)&=0\kern 5.0pt\text{and}\kern 5.0pt\widehat{\omega}(0)=0.\end{aligned}\right$.
@@
$\mathcal{T}(v,\pi,Q)\coloneqq\mathbb{O}(t)^{\top}\mathrm{T}(\mathbb{O}(t)v(t,y),\pi(t,y),Q(t,y))\mathbb{O}(t)$
for the rotation matrix $\mathbb{O}(t)$ related to (5.1), $\mathrm{T}(v,\pi,Q)$ as in (6.4) and $(z,\pi)=(\widehat{z}+z^{*},\widehat{\pi}+\pi^{*})$, we have
[…]
@@
[…]